OpenStudy (anonymous):

A rectangular sign had a perimeter of 44 feet and a length of 14 feet. A designer shortened the length so that the sign became a square with the same width as the old sign. What was the area of the new sign? A. 64 ft2 B. 196 ft2 C. 112 ft2 D. 256 ft2

OpenStudy (anonymous):

A rectangular sign had a perimeter of 44 feet and a length of 14 feet. A designer shortened the length so that the sign became a square with the same width as the old sign. What was the area of the new sign? A. 64 ft2 B. 196 ft2 C. 112 ft2 D. 256 ft2

OpenStudy (anonymous):

a length of 14 feet and a perimeter of 44 feet means that your width can be obtained with a perimeter function: perimeter is a function of length and width p(l, w) 2l + 2w = p(l, w) 2(14) + 2w = 44 = p(14, w) 28 + 2w = 44 2w = 44 - 28 2w = 16 w = 8 so if you have a new square with a width of 8 the value of its area is the function: a(w) = w^2 a(8) = (8)^2 a(8) = 64 so your area will be 64 ft squared